# A Mean-field Model of Optimal Investment

### v1

@@ 1. Introduction
In this paper, we consider a mean-field model of optimal investment with competition à la Cournot, where the price of the good produced by a representative company depends on the aggregate production of the entire economy through a nonlinear inverse demand function. In the absence of interventions, the production capacity of the representative company evolves stochastically over time as a geometric Brownian motion, and its level can be increased through investment, subject to quadratic costs. The representative company discounts profits and costs at a constant rate and aims to maximize total expected profits from production, net of investment costs. Instantaneous profits depend linearly on the company’s production capacity (thus, production occurs at full capacity) and on the time-dependent price of the produced good. The mean-field equilibrium investment and average production processes $(\widehat{\bf{u}},\widehat{q})$ are such that expected total net profits are maximized and, assuming an isoelastic inverse demand function, the price is given in terms of a nonlinear function of the average optimally controlled production at each time (see also Achdou et al. [2] and Remark 2.4 below). We are able to prove the existence and uniqueness of the equilibrium pair $(\widehat{\bf{u}},\widehat{q})$ when the problem’s time horizon is finite or infinite. Furthermore, we demonstrate that the existence and uniqueness carry over to the deterministic counterpart of our model, essentially using the same techniques as in the stochastic case.
@@
In our context, the consistency condition that the equilibrium price aligns at each time with a decreasing nonlinear function of the expected (optimally controlled) production capacity can be viewed as the limit, as the number $N$ of identical and indistinguishable companies operating in the market diverges, of the requirement that price inversely depends on the aggregate production of the entire economy, scaled by a factor of $1/N$. As discussed in Huang et al. [15], this scaling can be justified by considering situations where "an increasing number of firms join together to serve an increasing number of consumers" (see the discussion in Huang et al. [15], after Equation (2.4) therein).
@@
The equilibrium construction for a given time horizon $T\in(0,+\infty]$ follows a three-step approach. Firstly, given a deterministic path of the average production $q\coloneqq(q_{s})_{s\in[0,T]}$, we solve the representative company’s optimal investment problem. Because the production capacity evolves as a geometric Brownian motion and is linearly dependent on the investment process and since the performance criterion has linear dependence on the production capacity and quadratic dependence on the investment costs, the resulting optimal control problem for the representative company is of linear-quadratic type and has an explicit solution. Moreover, the optimal control $\widehat{\mathbf{u}}^{(T,q)}$ for fixed average production trajectory $q$ is deterministic. In the second step, we calculate the expectation of the optimally controlled production capacity process, which is easily computable given the explicit representation of the state process. Finally, in the third step, we impose the consistency condition, requiring that this expectation must match $q_{s}$ for each time $s\in[0,T]$. This condition leads to a nonlinear integral equation for the equilibrium average production trajectory $\widehat{q}\coloneqq(\widehat{q}_{s})_{s\in[0,T]}$. By construction the control $\widehat{\mathbf{u}}\coloneqq\widehat{\mathbf{u}}^{(T,\widehat{q})}$ and the function $\widehat{q}$ form a mean-field equilibrium.
@@
It is important to note that although the company’s optimal control problem for a fixed average production trajectory $q$ is linear-quadratic, due to the nonlinear dependence of the net profit functional on $q$, the overall mean-field problem is not of linear-quadratic type. In this regard, our results differ from those presented in Bensoussan et al. [3], Delarue and Tchuendom [12], Tchuendom [19], among others, which focus on linear-quadratic mean-field games.
@@
While the two-step approach previously outlined works for both the case where $T<+\infty$ and the case where $T=+\infty$, different technical arguments are used to prove the existence and uniqueness of the equilibrium in the finite and infinite horizon cases. As a matter of fact, the integral equation uniquely characterizing the equilibrium average production $\widehat{q}$ does not fall into the standard theory of integro-differential equations, as we have an initial-value problem with the integral being backward in time, and thus requires a careful ad hoc analysis. Specifically, the proof of the existence of a solution to it relies on a priori estimates and Schauder’s fixed point theorem when $T<+\infty$, whereas it utilizes a priori estimates and the Frechet-Kolmorogorov ${\mathrm{L}}^{p}$ compactness theorem in the case of $T=+\infty$. Uniqueness is then established in both cases through suitable contradiction arguments that exploit the properties of any mean-field equilibrium. Notably, when $T=+\infty$, we can also prove that the time-dependent equilibrium $\widehat{q}$ monotonically converges to the stationary average production level. The latter is obtained as the unique constant solution to the integral equation that uniquely determines $\widehat{q}$.
@@
To the best of our knowledge, the existence and uniqueness of the (nonstationary) mean-field equilibrium for a mean-field model of optimal investment with an isoelastic demand function, as discussed in this paper, is presented here for the first time.
The rest of the paper is organized as follows. The stochastic mean-field game is introduced in Section 2. In Section 3 the company’s optimal investment problem is solved for both the cases $T<+\infty$ and $T=+\infty$, while existence and uniqueness of the equilibrium is shown in Section 4, again for the finite and infinite time horizon cases. The deterministic version of the mean-field problem is finally considered in Section 5.
@@ 2. The stochastic model
We observe that the solution to SDE (2.1) can be written explicitly (cf. [16, Problem 5.6.15]), for each ${\mathbf{u}}\in{\mathcal{U}}_{T}$, as
@@
In our mean-field game we assume that every player aims at maximizing the discounted net profit functional
$J_{T,q}(\xi,{\mathbf{u}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho s}\left\{X_{s}^{\xi,{\mathbf{u}}}q_{s}^{-\beta}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right],$ (2.7)
where $\rho>0$ is a discount factor, $\beta>0$ is a fixed parameter, and $q=(q_{s})_{s\in[0,T]}$ is a given deterministic measurable function. At equilibrium, the function $q$ will identify with the average production capacity of the whole population of agents. This is formalized in the next definition.
@@
Remark 2.4.
The profit function appearing in (2.7) – i.e., the function $(x,p)\mapsto xp^{-\beta}$ – is related to the isoelastic demand obtained from Spence-Dixit-Stiglitz preferences and can be motivated as follows. [^1] See also the discussion in Achdou et al. [2, p. 7, Footnote 5] concerning the the model in Luttmer [18].
Assume that each player in our economy is a firm indexed by its productivity (or size) $x>0$. We can consider the firms’ production capacities as a proxy for this index and, thus, this setting is coherent with the model introduced above. Each firm produces a single good and faces the demand function
$\left(\frac{\pi(x)}{P}\right)^{-\gamma},\quad\text{with }\gamma>1,$
where $\pi(x)$ is the price set by a firm with productivity level $x$, and $P=\left[\int_{\mathbb{R}}\pi(x)^{1-\gamma}\mu({\mathrm{d}}x)\right]^{1/(1-\gamma)}$ is a price index, which is computed according to the statistical distribution $\mu$ of the firms productivity in the economy. Here we are assuming that the price set by two firms having an equal productivity level is the same. In other words, the price policy of each firm is exclusively determined by its productivity.
Each firm (i.e., fix $x>0$) aims at maximizing the function
$\pi\mapsto\left\{\pi\left(\frac{\pi}{P}\right)^{-\gamma}-\frac{1}{x}\left(\frac{\pi}{P}\right)^{-\gamma}\right\}.$
@@
$\frac{1}{\gamma-1}\left(\frac{\gamma}{\gamma-1}\right)^{-\gamma}x^{\gamma-1}P^{\gamma}\,,$ (2.8)
which is attained setting the price $\pi(x)=\frac{\gamma}{(\gamma-1)x}$.
This entails that $P=\frac{\gamma}{(\gamma-1)}\left(\int_{\mathbb{R}}x^{\gamma-1}\mu({\mathrm{d}}x)\right)^{1/(1-\gamma)}$, and hence, substituting this expression into (2.8), we get that the profit is
$\frac{1}{\gamma-1}x^{\gamma-1}p^{-\gamma},\quad\text{with}\quad p=\left(\int_{\mathbb{R}}x^{\gamma-1}\,\mu({\mathrm{d}}x)\right)^{\frac{1}{\gamma-1}}\,.$ (2.9)
Therefore, setting $\beta=\gamma=2$, we get the profit function $(x,p)\mapsto xp^{-\beta}$ appearing in our model. Note, however, that in the discounted net profit functional introduced in (2.7) we do not restrict the choice of the parameter $\beta$ to the one dictated by the economic application outlined above. Instead, we consider any possible value $\beta>0$, as the mathematical results that we prove can be stated in this more general framework.
The discussion above highlights that it is possible to consider more general forms of interaction between players in the profit function. In this paper we consider the average of the production capacities of the firms in the economy, but in (2.9) the geometric average appears. More generally, one can consider
$p=F\left(\int_{\mathbb{R}}f(x)\,\mu({\mathrm{d}}x)\right),$
for some $F,f\colon\mathbb{R}\to[0,+\infty)$ strictly increasing, see, e.g., [6]. However, as will be clear later on, the analysis carried out in this work heavily relies on the fact that the interaction between players is through the average of the production capacities. Indeed, thanks to this feature, we will find an explicit characterization of the solution to the mean-field game introduced above.

### v2

# Existence and Uniqueness Results for a Mean-field Game of Optimal Investment

@@ 1. Introduction
In this paper, we provide existence, uniqueness, and characterization results for the equilibrium of a mean-field model of optimal investment with competition à la Cournot. In the absence of interventions, the production capacity of the representative company evolves stochastically over time as a geometric Brownian motion, and its level can be increased through investment, subject to quadratic costs. The representative company discounts profits and costs at a constant rate and aims to maximize total expected profits from production, net of investment costs. Instantaneous profits depend linearly on the company’s production capacity (thus, production occurs at full capacity) and on the time-dependent price of the produced good. The mean-field equilibrium investment and average production processes $(\widehat{\bf{u}},\widehat{q})$ are such that expected total net profits are maximized and, assuming an isoelastic inverse demand function, the price is given in terms of a nonlinear function of the average optimally controlled production at each time (see also Achdou et al. [2] and Section 2.1 below). We are able to prove the existence and uniqueness of the equilibrium pair $(\widehat{\bf{u}},\widehat{q})$ when the problem’s time horizon is finite or infinite. These equilibria are characterized as the unique solution to a nonstandard system of forward-backward ordinary differential equations. Furthermore, we demonstrate that the existence and uniqueness carry over to the deterministic counterpart of our model, essentially using the same techniques as in the stochastic case.
@@
In our context, the consistency condition that the equilibrium price aligns at each time with a decreasing nonlinear function of the expected (optimally controlled) production capacity can be viewed as the limit, as the number $N$ of identical and indistinguishable companies operating in the market diverges, of the requirement that price inversely depends on the aggregate production of the entire economy, scaled by a factor of $1/N$. As discussed in Huang et al. [16], this scaling can be justified by considering situations where “an increasing number of firms join together to serve an increasing number of consumers” (see the discussion in Huang et al. [16], after Equation (2.4) therein).
@@
The equilibrium construction for a given time horizon $T\in(0,+\infty]$ follows a three-step approach. Firstly, given a deterministic path of the average production $q\coloneqq(q_{s})_{s\in[0,T]}$, we solve the representative company’s optimal investment problem. This is of linear-quadratic type and has an explicit solution, since the production capacity evolves as a geometric Brownian motion and is linearly dependent on the investment process, and because the performance criterion has linear dependence on the production capacity and quadratic dependence on the investment cost. Moreover, the optimal control $\widehat{\mathbf{u}}^{(T,q)}$ is deterministic, for any fixed average production trajectory $q$. In the second step, we calculate the expectation of the optimally controlled production capacity process, which is easily computable given the explicit representation of the state process. Finally, in the third step, we impose the consistency condition, requiring that this expectation must match $q_{s}$ for each time $s\in[0,T]$. This condition leads to a nonlinear integral equation for the equilibrium average production trajectory $\widehat{q}\coloneqq(\widehat{q}_{s})_{s\in[0,T]}$. By construction the control $\widehat{\mathbf{u}}\coloneqq\widehat{\mathbf{u}}^{(T,\widehat{q})}$ and the function $\widehat{q}$ form a mean-field equilibrium.
@@
It is important to note that although the company’s optimal control problem for a fixed average production trajectory $q$ is linear-quadratic, the overall mean-field problem is not of linear-quadratic type, due to the nonlinear dependence of the net profit functional on $q$. In this regard, our results differ from those presented in Bensoussan et al. [3], Delarue and Tchuendom [12], Tchuendom [20], among others, which focus on linear-quadratic mean-field games.
@@
While the three-step approach previously outlined works for the case where $T<+\infty$ as well as the case where $T=+\infty$, different technical arguments are used to prove the existence and uniqueness of the equilibrium in the finite and infinite horizon cases. As a matter of fact, the integral equation that uniquely characterizes the equilibrium average production $\widehat{q}$ does not fall into the standard theory of integro-differential equations, as we have an initial-value problem with the integral being backward in time, and this requires a careful ad hoc analysis. Specifically, the proof of the existence of a solution relies on a priori estimates and Schauder’s fixed point theorem when $T<+\infty$, whereas it utilizes a priori estimates and the Frechet-Kolmorogorov ${\mathrm{L}}^{p}$ compactness theorem in the case of $T=+\infty$. Uniqueness is then established in both cases through suitable contradiction arguments that exploit the properties of mean-field equilibria. Notably, when $T=+\infty$, we can also prove that the time-dependent equilibrium $\widehat{q}$ monotonically converges to the stationary average production level. The latter is obtained as the unique constant solution to the integral equation that uniquely determines $\widehat{q}$.
@@
To the best of our knowledge, the existence and uniqueness of the (nonstationary) mean-field equilibrium for a mean-field model of optimal investment with an isoelastic demand function, as discussed in this paper, is presented here for the first time. It is also worth noting that some of our results – in particular the existence and uniqueness ones – carry over also to the case of a more general instantaneous profit function and more general mean-field interaction, upon requiring suitable monotonicity conditions on the problem’s data. We refer to Section 2.1 below for further details.
The rest of the paper is organized as follows. The stochastic mean-field game is introduced in Section 2, and in Subsection 2.1 we provide its economic foundation. In Section 3 the company’s optimal investment problem is solved for both the cases $T<+\infty$ and $T=+\infty$, while existence and uniqueness of the equilibrium is shown in Section 4, again for the finite and infinite time horizon cases. The deterministic version of the mean-field problem is finally considered in Section 5.
@@ 2. The stochastic model
We observe that the solution to SDE (2.1) can be written explicitly (cf. [17, Problem 5.6.15]), for each ${\mathbf{u}}\in{\mathcal{U}}_{T}$, as
@@
In our mean-field game, the representative company’s instantaneous profit function is derived from the isoelastic demand function of Spence-Dixit-Stiglitz preferences (see Achdou et al. [2, p. 7, Footnote 5] and Luttmer [19]), as explained in Section 2.1 below. Moreover, the company faces a quadratic cost of investment. This leads to the expected net profit functional
$J_{T,q}(\xi,{\mathbf{u}})\coloneqq{\mathbb{E}}\left[\int_{0}^{T}{\mathrm{e}}^{-\rho s}\left\{X_{s}^{\xi,{\mathbf{u}}}q_{s}^{-\beta}-\dfrac{1}{2}u_{s}^{2}\right\}{\mathrm{d}}s\right]\,.$ (2.7)
Here, $\rho>0$ is a discount factor, $\beta>0$ is a fixed parameter, and $q=(q_{s})_{s\in[0,T]}$ is a given deterministic measurable function. At equilibrium, the function $q$ will identify with the average production capacity of the whole population of agents. This is formalized in the next definition.
@@
2.1. The economic foundation
To provide an economic foundation of our model, let us introduce the $N$-player game corresponding to the mean-field game introduced above.
We consider a finite number of firms $N$, whose production capacities evolve stochastically over time according to the following dynamics
[…]
where $B^{1},\dots,B^{n}$ are independent Wiener processes and $\xi^{1},\dots,\xi^{N}$ are i.i.d., ${\mathcal{F}}_{0}$-measurable, positive, and integrable random variables, independent from the Brownian motions $B^{i}$, $i=1,\dots,n$.
Each firm may increase its own production capacity, by investing at rate ${\mathbf{u}}^{i}=(u^{i}_{s})_{s\in[0,T]}\in{\mathcal{U}}_{T}$, and aims at maximizing the expected net profit functional
${\mathbb{E}}\Biggl{[}\int_{0}^{T}{\mathrm{e}}^{-\rho s}\Biggl{\{}X^{i}_{s}\biggl{(}\frac{1}{N}\sum_{j=1}^{N}X^{j}_{s}\biggr{)}^{-\beta}-\dfrac{1}{2}(u^{i}_{s})^{2}\Biggr{\}}{\mathrm{d}}s\Biggr{]}\,.$ (2.9)
Taking the limit as $N\to\infty$ and considering Nash equilibria, we expect that the dynamics (2.8) of each firm converges to the dynamics (2.1) of the representative player of the mean-field game, and that at equilibrium the functional (2.9) converges to $J_{T,\widehat{q}}$ given in (2.7), with $\widehat{q}$ satisfying the consistency condition of Definition 2.3-(ii).
We leave for future research the study of the $N$-player game and its convergence to the corresponding mean-field game, since this requires a detailed and careful analysis which is outside the scope of this paper, as we intend to focus on directly studying the mean-field game.
Now, instead, we want to focus on providing the precise derivation of the instantaneous profit function appearing in (2.9), i.e., the function
$(x_{1},\dots,x_{n})\mapsto x_{i}\biggl{(}\frac{1}{N}\sum_{j=1}^{N}x^{j}\biggr{)}^{-\beta},$ (2.10)
for any fixed $i=1,\dots,N$. This in turn justifies the form of the profit function $(x,p)\mapsto xp^{-\beta}$ appearing in (2.7). We follow an approach close to the one discussed in Achdou et al. [2, p. 7, Footnote 5].
Assume that each player $i=1,\dots,N$ in our economy is a firm producing a single good and facing the demand function
$\left(\frac{\pi_{i}}{P}\right)^{-\gamma},\quad\text{with }\gamma>1,$
where $\gamma$ is the elasticity of substitution, $\pi$ is the price set by firm $i$ and $P=\left(\frac{1}{N}\sum_{j=1}^{N}\pi_{j}^{1-\gamma}\right)^{1/(1-\gamma)}$ is a price index. This choice of the demand function is standard in economics and is justified by the Spence-Dixit-Stiglitz preferences (see, e.g., Dixit and Stiglitz [13]).
Each firm aims at maximizing the function
$\pi_{i}\mapsto\left\{\pi_{i}\left(\frac{\pi_{i}}{P}\right)^{-\gamma}-\frac{1}{x_{i}^{\delta}}\left(\frac{\pi_{i}}{P}\right)^{-\gamma}\right\},$
where $x_{i}>0$ is the productivity (or size) of firm $i$ and, hence, costs decrease with respect to productivity, according to the factor $x_{i}^{-\delta}$, where $\delta>0$ is a given parameter. We can consider each firm production capacity as a proxy for $x_{i}$ and, thus, this setting is coherent with the model introduced above.
@@ 2.1. The economic foundation
$\frac{1}{\gamma-1}\left(\frac{\gamma}{\gamma-1}\right)^{-\gamma}x_{i}^{\delta(\gamma-1)}P^{\gamma}\,,$ (2.11)
which is attained setting the price $\pi_{i}=\frac{\gamma}{(\gamma-1)}x_{i}^{-\delta}$.
This entails that $P=\frac{\gamma}{(\gamma-1)}\left(\frac{1}{N}\sum_{j=1}^{N}x_{j}^{\delta(\gamma-1)}\right)^{-1/(\gamma-1)}$, and hence, substituting this expression into (2.11), we get that the profit is
$\frac{1}{\gamma-1}x_{i}^{\delta(\gamma-1)}\left(\frac{1}{N}\sum_{j=1}^{N}x_{j}^{\delta(\gamma-1)}\right)^{-\frac{\gamma}{\gamma-1}}\,.$ (2.12)
Therefore, setting $\beta=\frac{\gamma}{\gamma-1}$, $\delta=\frac{1}{\gamma-1}$, and $p=\frac{1}{N}\sum_{j=1}^{N}x_{j}$, we get the profit function (2.10) appearing in our model, except for the constant $\frac{1}{\gamma-1}$ which does not affect the optimization problem.
The discussion above highlights that it is possible to consider more general forms of interaction between players in the profit function, which can be taken of the form
$\Phi(x)\,F\left(\frac{1}{N}\sum_{j=1}^{N}f(x_{j})\right),$ (2.13)
for some $\Phi,f\colon\mathbb{R}\to[0,+\infty)$ strictly increasing and $F\colon[0,+\infty)\to[0,+\infty)$, see, e.g., [6]. However, with the aim of obtaining an explicit characterization of the equilibrium, in this paper we focus on the particular case in which $\Phi$ and $f$ are the identity and $F(x)=x^{-\beta}$, as in (2.12). Nonetheless, it is worth noting that part of our subsequent results – in particular the existence ones – could be extended to the more general case of profit function (2.13), upon exploiting the monotonicity of functions $\Phi,F,$ and $f$.